# Yaohui Li2, Jun Lan1, Huijia Zhu1, Jianfu Zhang3, Weiqiang Wang1,Huaxiong Li2 1Ant Group

### v2

@@ 5.1 Implementation details
F3Net Image ACC 83.45 $77.55_{{\color[rgb]{0,0.453125,0}-5.90}}$ $82.04_{{\color[rgb]{0,0.453125,0}-1.41}}$ $81.45_{{\color[rgb]{0,0.453125,0}-2.00}}$ $69.71_{{\color[rgb]{0,0.453125,0}-13.74}}$ $76.90_{{\color[rgb]{0,0.453125,0}-7.55}}$ $78.67_{{\color[rgb]{0,0.453125,0}-4.78}}$ $82.94_{{\color[rgb]{0,0.453125,0}-0.51}}$ $82.93_{{\color[rgb]{0,0.453125,0}-0.52}}$
AP 88.73 $87.39_{{\color[rgb]{0,0.453125,0}-1.34}}$ $\textbf{81.02}_{{\color[rgb]{0,0.453125,0}-7.71}}$ $87.70_{{\color[rgb]{0,0.453125,0}-1.73}}$ $67.05_{{\color[rgb]{0,0.453125,0}-21.68}}$ $87.01_{{\color[rgb]{0,0.453125,0}-1.72}}$ $87.93_{{\color[rgb]{0,0.453125,0}-0.70}}$ $87.17_{{\color[rgb]{0,0.453125,0}-1.56}}$ $87.17_{{\color[rgb]{0,0.453125,0}-1.56}}$
NPR Image ACC 85.30 $80.83_{{\color[rgb]{0,0.453125,0}-4.47}}$ $73.92_{{\color[rgb]{0,0.453125,0}-11.38}}$ $83.82_{{\color[rgb]{0,0.453125,0}-1.48}}$ $71.69_{{\color[rgb]{0,0.453125,0}-13.61}}$ $83.26_{{\color[rgb]{0,0.453125,0}-2.04}}$ $82.43_{{\color[rgb]{0,0.453125,0}-2.87}}$ $84.94_{{\color[rgb]{0,0.453125,0}-0.36}}$ $85.03_{{\color[rgb]{0,0.453125,0}-0.47}}$
AP 82.45 $76.12_{{\color[rgb]{0,0.453125,0}-6.33}}$ $53.72_{{\color[rgb]{0,0.453125,0}-28.73}}$ $79.05_{{\color[rgb]{0,0.453125,0}-3.40}}$ $36.36_{{\color[rgb]{0,0.453125,0}-46.09}}$ $67.16_{{\color[rgb]{0,0.453125,0}-15.28}}$ $77.02_{{\color[rgb]{0,0.453125,0}-5.43}}$ $57.55_{{\color[rgb]{0,0.453125,0}-24.90}}$ $76.81_{{\color[rgb]{0,0.453125,0}-5.64}}$
STIL Video ACC 83.71 $78.14_{{\color[rgb]{0,0.453125,0}-5.57}}$ $73.79_{{\color[rgb]{0,0.453125,0}-9.92}}$ $81.97_{{\color[rgb]{0,0.453125,0}-1.74}}$ $77.58_{{\color[rgb]{0,0.453125,0}-6.13}}$ $77.64_{{\color[rgb]{0,0.453125,0}-6.07}}$ $74.16_{{\color[rgb]{0,0.453125,0}-9.55}}$ $83.44_{{\color[rgb]{0,0.453125,0}-0.27}}$ $83.47_{{\color[rgb]{0,0.453125,0}-0.24}}$
AP 87.12 $85.04_{{\color[rgb]{0,0.453125,0}-2.08}}$ $59.38_{{\color[rgb]{0,0.453125,0}-27.74}}$ $85.21_{{\color[rgb]{0,0.453125,0}-1.91}}$ $76.38_{{\color[rgb]{0,0.453125,0}-10.74}}$ $82.59_{{\color[rgb]{0,0.453125,0}-4.53}}$ $83.09_{{\color[rgb]{0,0.453125,0}-4.03}}$ $74.54_{{\color[rgb]{0,0.453125,0}-12.58}}$ $86.65_{{\color[rgb]{0,0.453125,0}-0.47}}$
CLIP-B-FT Image ACC 91.19 $82.92_{{\color[rgb]{0,0.453125,0}-8.27}}$ $76.78_{{\color[rgb]{0,0.453125,0}-12.41}}$ $89.23_{{\color[rgb]{0,0.453125,0}-12.41}}$ $66.91_{{\color[rgb]{0,0.453125,0}-24.28}}$ $79.65_{{\color[rgb]{0,0.453125,0}-11.54}}$ $78.58_{{\color[rgb]{0,0.453125,0}-12.61}}$ $88.50_{{\color[rgb]{0,0.453125,0}-2.69}}$ $88.81_{{\color[rgb]{0,0.453125,0}-2.38}}$
AP 91.52 $90.42_{{\color[rgb]{0,0.453125,0}-1.10}}$ $85.33_{{\color[rgb]{0,0.453125,0}-12.41}}$ $91.05_{{\color[rgb]{0,0.453125,0}-0.14}}$ $76.35_{{\color[rgb]{0,0.453125,0}-15.17}}$ $88.94_{{\color[rgb]{0,0.453125,0}-2.58}}$ $91.05_{{\color[rgb]{0,0.453125,0}-0.47}}$ $84.55_{{\color[rgb]{0,0.453125,0}-6.97}}$ $91.05_{{\color[rgb]{0,0.453125,0}-0.50}}$
XCLIP-B-FT Video ACC 85.66 $84.17_{{\color[rgb]{0,0.453125,0}-1.49}}$ $53.96_{{\color[rgb]{0,0.453125,0}-31.70}}$ $84.48_{{\color[rgb]{0,0.453125,0}-1.18}}$ $56.24_{{\color[rgb]{0,0.453125,0}-30.42}}$ $85.06_{{\color[rgb]{0,0.453125,0}-1.60}}$ $82.84_{{\color[rgb]{0,0.453125,0}-2.82}}$ $82.94_{{\color[rgb]{0,0.453125,0}-2.72}}$ $85.41_{{\color[rgb]{0,0.453125,0}-0.25}}$
AP 86.77 $86.53_{{\color[rgb]{0,0.453125,0}-0.24}}$ $54.11_{{\color[rgb]{0,0.453125,0}-32.66}}$ $86.76_{{\color[rgb]{0,0.453125,0}-0.01}}$ $46.10_{{\color[rgb]{0,0.453125,0}-40.67}}$ $86.71_{{\color[rgb]{0,0.453125,0}-0.06}}$ $86.74_{{\color[rgb]{0,0.453125,0}-0.03}}$ $68.05_{{\color[rgb]{0,0.453125,0}-14.89}}$ $86.01_{{\color[rgb]{0,0.453125,0}-0.76}}$
DeMamba Video ACC 92.29 $85.63_{{\color[rgb]{0,0.453125,0}-6.66}}$ $83.49_{{\color[rgb]{0,0.453125,0}-8.80}}$ $90.59_{{\color[rgb]{0,0.453125,0}-1.70}}$ $62.25_{{\color[rgb]{0,0.453125,0}-30.04}}$ $83.34_{{\color[rgb]{0,0.453125,0}-8.95}}$ $87.14_{{\color[rgb]{0,0.453125,0}-5.15}}$ $90.63_{{\color[rgb]{0,0.453125,0}-1.66}}$ $91.50_{{\color[rgb]{0,0.453125,0}-0.79}}$
-CLIP-FT AP 93.45 $90.93_{{\color[rgb]{0,0.453125,0}-2.52}}$ $74.03_{{\color[rgb]{0,0.453125,0}-19.42}}$ $93.44_{{\color[rgb]{0,0.453125,0}-0.01}}$ $61.12_{{\color[rgb]{0,0.453125,0}-32.33}}$ $90.08_{{\color[rgb]{0,0.453125,0}-3.37}}$ $93.35_{{\color[rgb]{0,0.453125,0}-0.10}}$ $82.99_{{\color[rgb]{0,0.453125,0}-10.46}}$ $93.39_{{\color[rgb]{0,0.453125,0}-0.06}}$
DeMamba Video ACC 94.42 $\textbf{90.52}_{{\color[rgb]{0,0.453125,0}-3.90}}$ $\textbf{93.26}_{{\color[rgb]{0,0.453125,0}-1.16}}$ $\textbf{94.26}_{{\color[rgb]{0,0.453125,0}-0.16}}$ $\textbf{72.98}_{{\color[rgb]{0,0.453125,0}-22.44}}$ $\textbf{89.59}_{{\color[rgb]{0,0.453125,0}-4.83}}$ $\textbf{91.06}_{{\color[rgb]{0,0.453125,0}-3.36}}$ $\textbf{94.36}_{{\color[rgb]{0,0.453125,0}-0.06}}$ $\textbf{93.82}_{{\color[rgb]{0,0.453125,0}-0.60}}$
-XCLIP-FT AP 97.10 $\textbf{96.00}_{{\color[rgb]{0,0.453125,0}-1.10}}$ $76.29_{{\color[rgb]{0,0.453125,0}-20.81}}$ $\textbf{96.76}_{{\color[rgb]{0,0.453125,0}-0.34}}$ $\textbf{67.18}_{{\color[rgb]{0,0.453125,0}-29.92}}$ $\textbf{94.30}_{{\color[rgb]{0,0.453125,0}-2.80}}$ $\textbf{97.07}_{{\color[rgb]{0,0.453125,0}-0.03}}$ $\textbf{85.05}_{{\color[rgb]{0,0.453125,0}-12.05}}$ $\textbf{97.05}_{{\color[rgb]{0,0.453125,0}-0.05}}$
@@
Model Metric Sora Morph Gen2 HotShot Lavie Show-1 Moon Crafter Model Wild Real Avg.
@@
w/o Demamba ACC 82.14 99.57 93.62 61.29 79.36 69.71 97.92 99.79 77.14 83.59 98.14 $85.66_{{\color[rgb]{0,0.453125,0}-8.76}}$
Block AP 64.42 99.73 96.78 70.98 90.35 77.28 97.34 99.84 82.01 88.97 - $86.77_{{\color[rgb]{0,0.453125,0}-10.43}}$
w/o global ACC 97.33 99.85 99.45 67.66 93.55 97.88 100.00 100.00 91.34 86.54 98.55 $93.83_{{\color[rgb]{0,0.453125,0}-0.59}}$
feat AP 87.32 99.35 99.22 82.44 97.99 99.42 99.42 99.85 97.32 94.98 - $95.73_{{\color[rgb]{0,0.453125,0}-1.37}}$
DeMamba ACC 98.21 100.00 99.86 65.43 94.86 98.86 100.00 100.00 92.86 89.09 99.42 94.42
-XCLIP-FT AP 93.32 100.00 99.97 85.55 98.97 99.60 99.98 100.00 97.77 95.75 - 97.10
@@
Model Zone Metric Sora Morph Gen2 HotShot Lavie Show-1 Moon Crafter Model Wild Real Avg.

### v3

@@ 5.1 Implementation details
F3Net Image R 0.8188 $0.7721_{{\color[rgb]{0,0.453125,0}-0.0467}}$ $0.8108_{{\color[rgb]{0,0.453125,0}-0.0080}}$ $0.8041_{{\color[rgb]{0,0.453125,0}-0.0147}}$ $0.6782_{{\color[rgb]{0,0.453125,0}-0.1406}}$ $0.7513_{{\color[rgb]{0,0.453125,0}-0.0675}}$ $0.7668_{{\color[rgb]{0,0.453125,0}-0.0520}}$ $0.8025_{{\color[rgb]{0,0.453125,0}-0.0163}}$ $0.8102_{{\color[rgb]{0,0.453125,0}-0.0086}}$
F1 0.8024 $0.7967_{{\color[rgb]{0,0.453125,0}-0.0057}}$ $0.6984_{{\color[rgb]{0,0.453125,0}-0.1040}}$ $0.7945_{{\color[rgb]{0,0.453125,0}-0.0079}}$ $0.5988_{{\color[rgb]{0,0.453125,0}-0.2036}}$ $0.7858_{{\color[rgb]{0,0.453125,0}-0.0166}}$ $0.7944_{{\color[rgb]{0,0.453125,0}-0.0080}}$ $0.7910_{{\color[rgb]{0,0.453125,0}-0.0114}}$ $0.7912_{{\color[rgb]{0,0.453125,0}-0.0112}}$
AP 0.8873 $0.8739_{{\color[rgb]{0,0.453125,0}-0.0134}}$ $0.8102_{{\color[rgb]{0,0.453125,0}-0.0771}}$ $0.8770_{{\color[rgb]{0,0.453125,0}-0.0173}}$ $0.6705_{{\color[rgb]{0,0.453125,0}-0.2168}}$ $0.8701_{{\color[rgb]{0,0.453125,0}-0.0172}}$ $0.8793_{{\color[rgb]{0,0.453125,0}-0.0070}}$ $0.8717_{{\color[rgb]{0,0.453125,0}-0.0156}}$ $0.8717_{{\color[rgb]{0,0.453125,0}-0.0156}}$
NPR Image R 0.8408 $0.8012_{{\color[rgb]{0,0.453125,0}-0.0396}}$ $0.7375_{{\color[rgb]{0,0.453125,0}-0.1033}}$ $0.8365_{{\color[rgb]{0,0.453125,0}-0.0043}}$ $0.7144_{{\color[rgb]{0,0.453125,0}-0.1264}}$ $0.8285_{{\color[rgb]{0,0.453125,0}-0.0123}}$ $0.8200_{{\color[rgb]{0,0.453125,0}-0.0208}}$ $0.8392_{{\color[rgb]{0,0.453125,0}-0.0016}}$ $0.8378_{{\color[rgb]{0,0.453125,0}-0.0030}}$
F1 0.7164 $0.6382_{{\color[rgb]{0,0.453125,0}-0.0782}}$ $0.4588_{{\color[rgb]{0,0.453125,0}-0.2576}}$ $0.6896_{{\color[rgb]{0,0.453125,0}-0.0268}}$ $0.1655_{{\color[rgb]{0,0.453125,0}-0.5509}}$ $0.5580_{{\color[rgb]{0,0.453125,0}-0.1584}}$ $0.6583_{{\color[rgb]{0,0.453125,0}-0.0581}}$ $0.4791_{{\color[rgb]{0,0.453125,0}-0.2373}}$ $0.6564_{{\color[rgb]{0,0.453125,0}-0.0600}}$
AP 0.8245 $0.7612_{{\color[rgb]{0,0.453125,0}-0.0633}}$ $0.5372_{{\color[rgb]{0,0.453125,0}-0.2873}}$ $0.7905_{{\color[rgb]{0,0.453125,0}-0.0340}}$ $0.3636_{{\color[rgb]{0,0.453125,0}-0.4609}}$ $0.6716_{{\color[rgb]{0,0.453125,0}-0.1528}}$ $0.7702_{{\color[rgb]{0,0.453125,0}-0.0543}}$ $0.5755_{{\color[rgb]{0,0.453125,0}-0.2490}}$ $0.7681_{{\color[rgb]{0,0.453125,0}-0.0564}}$
STIL Video R 0.8222 $0.7712_{{\color[rgb]{0,0.453125,0}-0.0510}}$ $0.7215_{{\color[rgb]{0,0.453125,0}-0.1007}}$ $0.8076_{{\color[rgb]{0,0.453125,0}-0.0146}}$ $0.7678_{{\color[rgb]{0,0.453125,0}-0.0544}}$ $0.7694_{{\color[rgb]{0,0.453125,0}-0.0528}}$ $0.7401_{{\color[rgb]{0,0.453125,0}-0.0821}}$ $0.8205_{{\color[rgb]{0,0.453125,0}-0.0017}}$ $0.8195_{{\color[rgb]{0,0.453125,0}-0.0027}}$
F1 0.7823 $0.7412_{{\color[rgb]{0,0.453125,0}-0.0411}}$ $0.5124_{{\color[rgb]{0,0.453125,0}-0.2699}}$ $0.7435_{{\color[rgb]{0,0.453125,0}-0.0388}}$ $0.6589_{{\color[rgb]{0,0.453125,0}-0.1234}}$ $0.7276_{{\color[rgb]{0,0.453125,0}-0.0547}}$ $0.7301_{{\color[rgb]{0,0.453125,0}-0.0522}}$ $0.6472_{{\color[rgb]{0,0.453125,0}-0.1351}}$ $0.7566_{{\color[rgb]{0,0.453125,0}-0.0257}}$
AP 0.8712 $0.8504_{{\color[rgb]{0,0.453125,0}-0.0208}}$ $0.5938_{{\color[rgb]{0,0.453125,0}-0.2774}}$ $0.8521_{{\color[rgb]{0,0.453125,0}-0.0191}}$ $0.7638_{{\color[rgb]{0,0.453125,0}-0.1074}}$ $0.8259_{{\color[rgb]{0,0.453125,0}-0.0453}}$ $0.8309_{{\color[rgb]{0,0.453125,0}-0.0403}}$ $0.7454_{{\color[rgb]{0,0.453125,0}-0.1258}}$ $0.8665_{{\color[rgb]{0,0.453125,0}-0.0047}}$
MINTIME-CLIP-B Video R 0.8724 $0.8215_{{\color[rgb]{0,0.453125,0}-0.0509}}$ $0.8199_{{\color[rgb]{0,0.453125,0}-0.0525}}$ $0.8215_{{\color[rgb]{0,0.453125,0}-0.0508}}$ $0.5534_{{\color[rgb]{0,0.453125,0}-0.3190}}$ $0.8084_{{\color[rgb]{0,0.453125,0}-0.0640}}$ $0.7972_{{\color[rgb]{0,0.453125,0}-0.0752}}$ $0.8628_{{\color[rgb]{0,0.453125,0}-0.0096}}$ $0.8715_{{\color[rgb]{0,0.453125,0}-0.0009}}$
F1 0.8700 $0.8486_{{\color[rgb]{0,0.453125,0}-0.0214}}$ $\textbf{0.8417}_{{\color[rgb]{0,0.453125,0}-0.0283}}$ $0.8486_{{\color[rgb]{0,0.453125,0}-0.0214}}$ $0.5485_{{\color[rgb]{0,0.453125,0}-0.3215}}$ $0.8468_{{\color[rgb]{0,0.453125,0}-0.0232}}$ $0.8379_{{\color[rgb]{0,0.453125,0}-0.0321}}$ $0.4734_{{\color[rgb]{0,0.453125,0}-0.3966}}$ $0.8658_{{\color[rgb]{0,0.453125,0}-0.0042}}$
AP 0.9369 $0.9207_{{\color[rgb]{0,0.453125,0}-0.0158}}$ $\textbf{0.9143}_{{\color[rgb]{0,0.453125,0}-0.0226}}$ $0.9207_{{\color[rgb]{0,0.453125,0}-0.0158}}$ $0.6199_{{\color[rgb]{0,0.453125,0}-0.3170}}$ $0.9071_{{\color[rgb]{0,0.453125,0}-0.0298}}$ $0.9227_{{\color[rgb]{0,0.453125,0}-0.0142}}$ $0.8146_{{\color[rgb]{0,0.453125,0}-0.1223}}$ $0.9364_{{\color[rgb]{0,0.453125,0}-0.0005}}$
FTCN-CLIP-B Video R 0.8935 $0.8519_{{\color[rgb]{0,0.453125,0}-0.0416}}$ $0.8127_{{\color[rgb]{0,0.453125,0}-0.0808}}$ $0.8858_{{\color[rgb]{0,0.453125,0}-0.0077}}$ $0.5546_{{\color[rgb]{0,0.453125,0}-0.3389}}$ $0.8284_{{\color[rgb]{0,0.453125,0}-0.0651}}$ $0.8547_{{\color[rgb]{0,0.453125,0}-0.0388}}$ $0.8874_{{\color[rgb]{0,0.453125,0}-0.0061}}$ $0.8904_{{\color[rgb]{0,0.453125,0}-0.0031}}$
F1 0.8739 $0.8730_{{\color[rgb]{0,0.453125,0}-0.0009}}$ $0.5495_{{\color[rgb]{0,0.453125,0}-0.3244}}$ $0.8484_{{\color[rgb]{0,0.453125,0}-0.0451}}$ $0.5573_{{\color[rgb]{0,0.453125,0}-0.3166}}$ $0.8689_{{\color[rgb]{0,0.453125,0}-0.0050}}$ $0.8648_{{\color[rgb]{0,0.453125,0}-0.0091}}$ $0.6002_{{\color[rgb]{0,0.453125,0}-0.2735}}$ $0.8638_{{\color[rgb]{0,0.453125,0}-0.0101}}$
AP 0.9418 $0.9294_{{\color[rgb]{0,0.453125,0}-0.0124}}$ $0.6884_{{\color[rgb]{0,0.453125,0}-0.2534}}$ $0.9316_{{\color[rgb]{0,0.453125,0}-0.0102}}$ $0.6015_{{\color[rgb]{0,0.453125,0}-0.3403}}$ $0.9256_{{\color[rgb]{0,0.453125,0}-0.0162}}$ $0.9386_{{\color[rgb]{0,0.453125,0}-0.0032}}$ $0.8503_{{\color[rgb]{0,0.453125,0}-0.0915}}$ $0.9389_{{\color[rgb]{0,0.453125,0}-0.0029}}$
TALL Video R 0.8852 $0.8413_{{\color[rgb]{0,0.453125,0}-0.0439}}$ $0.4272_{{\color[rgb]{0,0.453125,0}-0.4580}}$ $0.8785_{{\color[rgb]{0,0.453125,0}-0.0067}}$ $0.7233_{{\color[rgb]{0,0.453125,0}-0.1619}}$ $0.8061_{{\color[rgb]{0,0.453125,0}-0.0791}}$ $0.8045_{{\color[rgb]{0,0.453125,0}-0.0807}}$ $0.8834_{{\color[rgb]{0,0.453125,0}-0.0018}}$ $0.8816_{{\color[rgb]{0,0.453125,0}-0.0036}}$
F1 0.7419 $0.7411_{{\color[rgb]{0,0.453125,0}-0.0008}}$ $0.5309_{{\color[rgb]{0,0.453125,0}-0.2110}}$ $0.7248_{{\color[rgb]{0,0.453125,0}-0.0171}}$ $0.5844_{{\color[rgb]{0,0.453125,0}-0.1575}}$ $0.7329_{{\color[rgb]{0,0.453125,0}-0.0090}}$ $0.7360_{{\color[rgb]{0,0.453125,0}-0.0059}}$ $0.5996_{{\color[rgb]{0,0.453125,0}-0.1423}}$ $0.6719_{{\color[rgb]{0,0.453125,0}-0.0700}}$
AP 0.8791 $0.8673_{{\color[rgb]{0,0.453125,0}-0.0118}}$ $0.6439_{{\color[rgb]{0,0.453125,0}-0.2352}}$ $0.8774_{{\color[rgb]{0,0.453125,0}-0.0017}}$ $0.6744_{{\color[rgb]{0,0.453125,0}-0.2047}}$ $0.8608_{{\color[rgb]{0,0.453125,0}-0.0183}}$ $0.8767_{{\color[rgb]{0,0.453125,0}-0.0024}}$ $0.7880_{{\color[rgb]{0,0.453125,0}-0.0911}}$ $0.8439_{{\color[rgb]{0,0.453125,0}-0.0352}}$
CLIP-B-FT Image R 0.9057 $0.8253_{{\color[rgb]{0,0.453125,0}-0.0804}}$ $0.7632_{{\color[rgb]{0,0.453125,0}-0.1425}}$ $0.8878_{{\color[rgb]{0,0.453125,0}-0.0179}}$ $0.6782_{{\color[rgb]{0,0.453125,0}-0.2275}}$ $0.7877_{{\color[rgb]{0,0.453125,0}-0.1180}}$ $0.7832_{{\color[rgb]{0,0.453125,0}-0.1225}}$ $0.8812_{{\color[rgb]{0,0.453125,0}-0.0245}}$ $0.8809_{{\color[rgb]{0,0.453125,0}-0.0248}}$
F1 0.7631 $0.7476_{{\color[rgb]{0,0.453125,0}-0.0155}}$ $0.6982_{{\color[rgb]{0,0.453125,0}-0.0649}}$ $0.7533_{{\color[rgb]{0,0.453125,0}-0.0098}}$ $0.6273_{{\color[rgb]{0,0.453125,0}-0.1358}}$ $0.7125_{{\color[rgb]{0,0.453125,0}-0.0506}}$ $0.7537_{{\color[rgb]{0,0.453125,0}-0.0094}}$ $0.6743_{{\color[rgb]{0,0.453125,0}-0.0888}}$ $0.7542_{{\color[rgb]{0,0.453125,0}-0.0089}}$
AP 0.9152 $0.9042_{{\color[rgb]{0,0.453125,0}-0.0110}}$ $0.8533_{{\color[rgb]{0,0.453125,0}-0.0619}}$ $0.9105_{{\color[rgb]{0,0.453125,0}-0.0047}}$ $0.7635_{{\color[rgb]{0,0.453125,0}-0.1517}}$ $0.8894_{{\color[rgb]{0,0.453125,0}-0.0258}}$ $0.9105_{{\color[rgb]{0,0.453125,0}-0.0047}}$ $0.8455_{{\color[rgb]{0,0.453125,0}-0.0697}}$ $0.9105_{{\color[rgb]{0,0.453125,0}-0.0047}}$
DeMamba-CLIP-FT Video R 0.9158 $0.8572_{{\color[rgb]{0,0.453125,0}-0.0586}}$ $0.8479_{{\color[rgb]{0,0.453125,0}-0.0679}}$ $0.9047_{{\color[rgb]{0,0.453125,0}-0.0111}}$ $0.6172_{{\color[rgb]{0,0.453125,0}-0.2986}}$ $0.8352_{{\color[rgb]{0,0.453125,0}-0.0806}}$ $0.8664_{{\color[rgb]{0,0.453125,0}-0.0494}}$ $0.9037_{{\color[rgb]{0,0.453125,0}-0.0121}}$ $0.9105_{{\color[rgb]{0,0.453125,0}-0.0053}}$
F1 0.8919 $0.8750_{{\color[rgb]{0,0.453125,0}-0.0169}}$ $0.6325_{{\color[rgb]{0,0.453125,0}-0.2594}}$ $0.8912_{{\color[rgb]{0,0.453125,0}-0.0007}}$ $0.5233_{{\color[rgb]{0,0.453125,0}-0.3686}}$ $0.8455_{{\color[rgb]{0,0.453125,0}-0.0464}}$ $0.8895_{{\color[rgb]{0,0.453125,0}-0.0024}}$ $0.7672_{{\color[rgb]{0,0.453125,0}-0.1247}}$ $0.8910_{{\color[rgb]{0,0.453125,0}-0.0009}}$
AP 0.9345 $0.9093_{{\color[rgb]{0,0.453125,0}-0.0252}}$ $0.7403_{{\color[rgb]{0,0.453125,0}-0.1942}}$ $0.9344_{{\color[rgb]{0,0.453125,0}-0.0001}}$ $0.6112_{{\color[rgb]{0,0.453125,0}-0.3233}}$ $0.9008_{{\color[rgb]{0,0.453125,0}-337}}$ $0.9335_{{\color[rgb]{0,0.453125,0}-0.0010}}$ $0.8299_{{\color[rgb]{0,0.453125,0}-0.1046}}$ $0.9339_{{\color[rgb]{0,0.453125,0}-0.0006}}$
XCLIP-B-FT Video R 0.8441 $0.8387_{{\color[rgb]{0,0.453125,0}-0.0054}}$ $0.5377_{{\color[rgb]{0,0.453125,0}-0.3064}}$ $0.8378_{{\color[rgb]{0,0.453125,0}-0.0063}}$ $0.5597_{{\color[rgb]{0,0.453125,0}-0.2844}}$ $0.8388_{{\color[rgb]{0,0.453125,0}-0.0023}}$ $0.8254_{{\color[rgb]{0,0.453125,0}-0.0187}}$ $0.8195_{{\color[rgb]{0,0.453125,0}-0.0216}}$ $0.8421_{{\color[rgb]{0,0.453125,0}-0.0020}}$
F1 0.7662 $0.7329_{{\color[rgb]{0,0.453125,0}-0.0333}}$ $0.4188_{{\color[rgb]{0,0.453125,0}-0.3474}}$ $0.7592_{{\color[rgb]{0,0.453125,0}-0.0070}}$ $0.3062_{{\color[rgb]{0,0.453125,0}-0.4600}}$ $0.7595_{{\color[rgb]{0,0.453125,0}-0.0067}}$ $0.7593_{{\color[rgb]{0,0.453125,0}-0.0069}}$ $0.6322_{{\color[rgb]{0,0.453125,0}-0.1340}}$ $0.7488_{{\color[rgb]{0,0.453125,0}-0.0174}}$
AP 0.8677 $0.8653_{{\color[rgb]{0,0.453125,0}-0.0024}}$ $0.5411_{{\color[rgb]{0,0.453125,0}-0.3266}}$ $0.8676_{{\color[rgb]{0,0.453125,0}-0.0001}}$ $0.4610_{{\color[rgb]{0,0.453125,0}-0.4067}}$ $0.8671_{{\color[rgb]{0,0.453125,0}-0.0006}}$ $0.8674_{{\color[rgb]{0,0.453125,0}-0.0003}}$ $0.6805_{{\color[rgb]{0,0.453125,0}-0.1489}}$ $0.8601_{{\color[rgb]{0,0.453125,0}-0.0076}}$
DeMamba-XCLIP-FT Video R 0.9392 $\textbf{0.9052}_{{\color[rgb]{0,0.453125,0}-0.0340}}$ $\textbf{0.9326}_{{\color[rgb]{0,0.453125,0}-0.0116}}$ $\textbf{0.9376}_{{\color[rgb]{0,0.453125,0}-0.0016}}$ $\textbf{0.7298}_{{\color[rgb]{0,0.453125,0}-0.2244}}$ $\textbf{0.8915}_{{\color[rgb]{0,0.453125,0}-0.0477}}$ $\textbf{0.9094}_{{\color[rgb]{0,0.453125,0}-0.0288}}$ $\textbf{0.9382}_{{\color[rgb]{0,0.453125,0}-0.0010}}$ $\textbf{0.9342}_{{\color[rgb]{0,0.453125,0}-0.0050}}$
F1 0.9020 $\textbf{0.8770}_{{\color[rgb]{0,0.453125,0}-0.0250}}$ $0.6823_{{\color[rgb]{0,0.453125,0}-0.2197}}$ $\textbf{0.8820}_{{\color[rgb]{0,0.453125,0}-0.0200}}$ $\textbf{0.6595}_{{\color[rgb]{0,0.453125,0}-0.2425}}$ $\textbf{0.8755}_{{\color[rgb]{0,0.453125,0}-0.0265}}$ $\textbf{0.8990}_{{\color[rgb]{0,0.453125,0}-0.0030}}$ $\textbf{0.7962}_{{\color[rgb]{0,0.453125,0}-0.1058}}$ $\textbf{0.9005}_{{\color[rgb]{0,0.453125,0}-0.0005}}$
AP 0.9710 $\textbf{0.9600}_{{\color[rgb]{0,0.453125,0}-0.0110}}$ $0.7629_{{\color[rgb]{0,0.453125,0}-0.2081}}$ $\textbf{0.9676}_{{\color[rgb]{0,0.453125,0}-0.0034}}$ $\textbf{0.6718}_{{\color[rgb]{0,0.453125,0}-0.2992}}$ $\textbf{0.9430}_{{\color[rgb]{0,0.453125,0}-0.0280}}$ $\textbf{0.9707}_{{\color[rgb]{0,0.453125,0}-0.0003}}$ $\textbf{0.8505}_{{\color[rgb]{0,0.453125,0}-0.1205}}$ $\textbf{0.9705}_{{\color[rgb]{0,0.453125,0}-0.0005}}$
@@
Model Metric Sora Morph Gen2 HotShot Lavie Show-1 Moon Crafter Model Wild Avg.
@@
w/o Demamba R 0.8214 0.9957 0.9362 0.6129 0.7936 0.6971 0.9792 0.9979 0.7714 0.8359 $0.8411_{{\color[rgb]{0,0.453125,0}-0.0971}}$
F1 0.3147 0.8805 0.9041 0.6530 0.8242 0.7099 0.8602 0.9370 0.7570 0.8212 $0.7662_{{\color[rgb]{0,0.453125,0}-0.1358}}$
AP 0.6442 0.9973 0.9678 0.7098 0.9035 0.7728 0.9734 0.9984 0.8201 0.8897 $0.8677_{{\color[rgb]{0,0.453125,0}-0.1043}}$
w/o global R 0.9733 0.9985 0.9945 0.6766 0.9355 0.9788 1.0000 1.0000 0.9134 0.8654 $0.9336_{{\color[rgb]{0,0.453125,0}-0.0046}}$
F1 0.4252 0.9053 0.9475 0.7183 0.9176 0.8960 0.8962 0.9508 0.8618 0.8562 $0.8375_{{\color[rgb]{0,0.453125,0}-0.0645}}$
AP 0.8732 0.9935 0.9922 0.8244 0.9799 0.9942 0.9942 0.9985 0.9732 0.9498 $0.9573_{{\color[rgb]{0,0.453125,0}-0.0137}}$
DeMamba-XCLIP-FT R 0.9821 1.0000 0.9986 0.6543 0.9486 0.9886 1.0000 1.0000 0.9286 0.8909 0.9382
F1 0.6467 0.9602 0.9790 0.7539 0.9537 0.9551 0.9557 0.9797 0.9240 0.9120 0.9020
AP 0.9332 1.0000 0.9997 0.8555 0.9897 0.9960 0.9998 1.0000 0.9777 0.9575 0.9710
@@
Model Zone Metric Sora Morph Gen2 HotShot Lavie Show-1 Moon Crafter Model Wild Avg.